# Three-dimensional Imaging of Pion using Lattice QCD: Generalized Parton Distributions

### v1

@@ Abstract
I Introduction
@@
Since its discovery in 1947, the pion has been the subject of intense research, recognized for its dual identity as both a Goldstone boson linked to chiral symmetry breaking and a QCD-bound state. Over the years, substantial efforts have been dedicated to studying its internal structure by analyzing experimental data. Key methodologies have involved extracting form factors (FFs) through the pion-electron scattering [1] and discerning parton distribution functions (PDFs) via the Drell-Yan process [2]. However, these approaches are limited to revealing only the one-dimensional structure of the hadron. For a more comprehensive, three-dimensional perspective, the focus has shifted to generalized parton distributions (GPDs), a concept introduced in the 1990s [3, 4, 5, 6]. Access to GPDs is typically gained through exclusive reactions, notably deeply virtual Compton scattering [6] and deeply virtual meson production [7, 8].
@@
Presently, a series of experiments are being conducted or planned, which can enrich our understanding of the pion internal structure. Prominent among these are the JLab 12 GeV program [9], the Apparatus for Meson and Baryon Experimental Research (AMBER) at CERN SPS [10], the forthcoming Electron-Ion Collider (EIC) at Brookhaven National Laboratory [11], and the Electron-Ion Collider in China (EicC) [12]. These initiatives are designed to probe the pion structure in various kinematic regimes. However, extracting GPDs from experimental data is fraught with challenges, including the chiral-odd nature of certain distributions, such as the transversity GPDs, and the complexity involved in pion production. On the other hand, lattice QCD results, offering complementary insights and potentially guiding experimental efforts, are highly sought after. The advent of Large Momentum Effective Theory (LaMET) [13, 14, 15] in 2013 made it possible to directly calculate the Bjorken-$x$ dependence of the parton distributions, extending the scope of GPD studies beyond just the first few Mellin moments [16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. This spurred a flurry of lattice research on nucleon and meson GPDs [26, 27, 28, 29, 30, 31, 32, 33, 34] based on the quasi-distribution (see reviews in Refs. [15, 35]), which is defined in terms of matrix elements of non-local operators involving quark and antiquark fields separated by a spatial distance, rather than a light-cone distance, and thus is calculable on the lattice.
@@
While there are several lattice QCD studies of the nucleon GPDs [27, 28, 29, 30, 31, 33, 34], including a few based on alternate methods [36, 37, 38, 39], there have been limited LaMET calculations of the pion GPDs [26, 32]. Traditionally, GPD calculations are conducted in the Breit (symmetric) frame, requiring the momentum transfer to be symmetrically distributed between the initial and final states. However, this approach incurs substantial computational costs. Recently, a frame-independent method was proposed to extract GPDs from lattice calculations in an arbitrary frame [30]. This innovative approach holds the potential to significantly reduce computational expenses by working in a non-Breit (asymmetric) frame where the initial or final state momentum is fixed.
@@
Two crucial aspects of obtaining the light-cone parton distributions from the lattice calculations are the renormalization and matching process. The GPDs and quasi-GPDs (qGPDs) are typically defined in different renormalization schemes. For the qGPDs extracted from the lattice analyses, renormalization is required to eliminate the ultraviolet (UV) divergences stemming from the Wilson line. Commonly employed renormalization methods include the regularization-independent momentum-subtraction (RI/MOM) [40, 41, 42, 43] and various ratio schemes [44, 45, 46, 47, 48]. However, these methods adhere to a factorization relation with the light-cone correlation only at short distances. At long distances, they introduce nonperturbative effects [49], which will impact the qGPDs through the Fourier transform of the matrix elements, thereby affecting the LaMET matching results in Bjorken-$x$ space. To overcome this issue, in this work, we use a hybrid-scheme renormalization [49]. The key point of this scheme is to renormalize the matrix elements at short and long distances separately, which can remove the linear divergence at long distances without introducing additional nonperturbative effects.
@@
Furthermore, we match the qGPDs in the lattice renormalization scheme to the light-cone GPDs in the $\overline{\rm MS}$ scheme through LaMET [50, 51, 52, 53]. During this process, the accuracy of the perturbative calculation plays a significant role in the precision of the final GPD results. In the zero-skewness limit, the matching kernel of GPDs is the same as the one for PDFs [51], which has been derived up to the next-to-next-to-leading order (NNLO) [47, 54]. We also utilize renormalization group resummation (RGR) [55, 56] to resum the small-$x$ logarithmic terms. Additionally, we consider leading-renormalon resummation (LRR) [57, 58], which can remove the renormalon ambiguity in the Wilson-line mass matching, to eliminate the linear power corrections [58]. Therefore, we achieve a state-of-the-art calculation of the valence pion GPDs using an adapted hybrid-scheme renormalization, along with the implementation of the combined NNLO+RGR+LRR matching.
@@
In this study, we present our lattice calculations of the valence pion GPDs using the LaMET approach, featuring a very small lattice spacing of 0.04 fm within the non-Breit frame. The organization of this paper is as follows: In Section II, we review the definitions of pion GPDs and qGPDs in an arbitrary frame, outline our lattice setup, and describe the methodologies employed to extract the matrix elements from the combined analyses of the two-point and three-point correlation functions. In this section, we also detail the renormalization procedure of the qGPD matrix elements in the coordinate space using the hybrid scheme and discuss the challenges of performing the Fourier transform to get the qGPDs. Subsequently, we apply the LaMET matching approach to derive the valence light-cone GPDs in Section III. This section contains our main results, including the sensitivity of our results to the perturbative accuracy of the matching and the dependence on the renormalization scales. Finally, Section IV provides a summary of our findings.
II Valence pion quasi-GPDs in asymmetric frame on the lattice
II.1 General considerations
@@
Our goal is to calculate the valence pion GPDs in an asymmetric frame using the frame-independent approach laid out in Ref. [30]. For this, we consider the following non-local iso-vector matrix element
@@
where ${\cal W}_{-\frac{z}{2},\frac{z}{2}}$ is the Wilson line that connects the quark and antiquark fields, ensuring the gauge invariant of the matrix element. Here, we use the following normalization of the pion states $\left\langle\pi(P)|\pi(P^{\prime})\right\rangle=E_{p}(2\pi)^{3}\delta^{3}(\mathbf{P}-\mathbf{P^{\prime}})$. Following Ref. [30], we write the matrix element in terms of the Lorentz-invariant amplitudes $A_{i}(z\cdot\bar{P},z\cdot\Delta,\Delta^{2},z^{2})$ for $i=1,2,3$, as follows:
@@
where the skewness parameter $\xi=-\Delta^{+}/(2P^{+})$ and $t=\Delta^{2}$. However, it is possible to define the light-cone GPDs in a frame-independent, i.e., Lorentz-invariant way [30] as
@@
In the Lorentz-invariant framework, the momentum-space GPD is defined as [30]
@@
$\widetilde{H}(x,\xi,t)=\int\frac{d(z\cdot\bar{P})}{2\pi}e^{ixz\cdot\bar{P}}\widetilde{H}_{\rm LI}(z\cdot\bar{P},-2\xi(z\cdot\bar{P}),t,z^{2}).$ (9)
By studying the behavior under hermiticity and time-reversal transformations simultaneously, it was found that the amplitude $A_{3}$ is an odd function of $z\cdot\Delta=-2\xi(z\cdot\bar{P})$ [30, 33]. In the forward limit, GPDs should smoothly approach PDFs, implying that $A_{3}$ should be equal to 0 in this work at zero skewness. In the following subsections, we will discuss our lattice setup and lattice QCD results on $A_{i}$.
II.2 Lattice setup
Our lattice calculations use the gauge ensembles provided by the HotQCD collaboration [59], utilizing a 2+1 flavor setup with Highly Improved Staggered Quark (HISQ) action [60]. The lattice configuration has dimensions of $N_{s}\times N_{t}=64^{3}\times 64$ and a lattice spacing of $a=0.04$ fm. The sea quark masses are adjusted to yield a pion mass of 160 MeV. In the valence sector, we use the Wilson-Clover action with one level of hypercubic (HYP) smearing [61]. For the coefficient of the clover term, we utilize the tree-level tadpole-improved value computed with the fourth root of the plaquette, yielding 1.02868 [62]. The valence quark masses in the Wilson-Clover action are tuned to $am_{q}=-0.033$, resulting in a pion mass of 300 MeV.
Central to our computational approach is the use of momentum-boosted smeared sources [63]. The quark propagators are obtained through the application of the multigrid algorithm [64] to invert the Wilson-Dirac operator using the QUDA software suite [65, 66, 67] on GPUs. In this work, we consider the pion boosted along the $z$-direction with momenta $P_{z}=2\pi n_{z}/(aN_{s})$ with $n_{z}$ being an integer. To obtain a good signal for the boosted pions, momentum-boosted sources and sinks are constructed using a Gaussian profile, with boost momenta $k_{z}=2\pi m_{z}/(aN_{s})$ in the $z$-direction. Source construction is done in the Coulomb gauge, and the Gaussian profile is created with the radius of 0.208 fm [68, 62]. Employing these quark propagators, we have computed both the two-point and three-point hadron correlation functions. To increase statistics per configuration, we combined multiple exact (high-precision) and sloppy (low-precision) samples and implemented the all-mode averaging (AMA) technique [69]. The lattice parameters used in this study are summarized in Table 1.
@@ II.2 Lattice setup
The approach used to obtain the pion matrix element here closely follows our previous works [62, 70, 71, 72, 73, 74]. We fit the large $t_{s}$ behavior of the two-point correlation functions with two or three states and obtain the energies of these states for different momenta. The corresponding energy levels are then used in the analysis of the three-point function and the extraction of the pion matrix elements $M^{\mu}$. We consider the following ratio of
$R^{\mu}(\mathbf{P}^{f},\mathbf{P}^{i};t_{s},\tau;z)\equiv\sqrt{E_{0}^{f}E_{0}^{i}}\frac{C_{{\rm 3pt}}^{\mu}(\mathbf{P}^{f},\mathbf{P^{i}};t_{s},\tau;z)}{C_{{\rm 2pt}}(\mathbf{P}^{f},t_{s})}\times\left[\frac{C_{{\rm 2pt}}(\mathbf{P}^{i},t_{s}-\tau)C_{{\rm 2pt}}(\mathbf{P}^{f},\tau)C_{{\rm 2pt}}(\mathbf{P}^{f},t_{s})}{C_{{\rm 2pt}}(\mathbf{P}^{f},t_{s}-\tau)C_{{\rm 2pt}}(\mathbf{P}^{i},\tau)C_{{\rm 2pt}}(\mathbf{P}^{i},t_{s})}\right]^{1/2},$ (12)
where $E_{0}^{i}$ and $E_{0}^{f}$ correspond to the ground-state pion energies in the initial and final states, respectively, and the kinematic factor $\sqrt{E_{0}^{f}E_{0}^{i}}$ is used for the normalization. For large $t_{s}$ and $\tau$, this ratio gives the matrix element $M^{\mu}$. We assume that for large $t_{s}$ and $\tau$, two or three states mainly contribute to $R^{\mu}(\mathbf{P}^{f},\mathbf{P}^{i};t_{s},\tau;z)$ and perform the corresponding fits to obtain the matrix element. Some details of the fits are provided in Appendix A.
Frame $t_{s}/a$ $\mathbf{n}^{f}=(n^{f}_{x},n^{f}_{y},n^{f}_{z})$ $m_{z}$ $P_{z}$[GeV] $\mathbf{n}^{\Delta}=(n^{\Delta}_{x},n^{\Delta}_{y},n^{\Delta}_{z})$ $-t$[GeV2] #cfgs (#ex, #sl)
Breit 9,12,15,18 (1, 0, 2) 2 0.968 (2, 0, 0) 0.938 115 (1, 32)
non-Breit 9,12,15,18 (0,0,0) 0 0 (0,0,0) 0 314 (3, 96)
9,12,15,18 (0,0,2) 2 0.968 (1,2,0) 0.952 314 (4, 128)
II.3 Lattice results on the Lorentz-invariant amplitudes

### v2

@@ Abstract
Keywords:
Lattice Quantum Chromodynamics, Pion, Generalized Parton Distributions
arxiv: 2407.03516
1 Introduction
@@
Since its discovery in 1947, the pion has been the subject of intense research, recognized for its dual identity as both a Goldstone boson linked to chiral symmetry breaking and a bound state in quantum chromodynamics (QCD). Over the years, substantial efforts have been dedicated to studying its internal structure by analyzing experimental data. Key methodologies have involved extracting form factors (FFs) through the pion-electron scattering NA7:1986vav and discerning parton distribution functions (PDFs) via the Drell-Yan process Conway:1989fs . However, these approaches are limited to revealing only the one-dimensional structure of the hadron. For a more comprehensive, three-dimensional perspective, the focus has shifted to generalized parton distributions (GPDs), a concept introduced in the 1990s Muller:1994ses ; Ji:1996ek ; Radyushkin:1996nd ; Ji:1996nm . Access to GPDs is typically gained through exclusive reactions, notably deeply virtual Compton scattering Ji:1996nm and deeply virtual meson production Radyushkin:1996ru ; Collins:1996fb .
@@
Presently, a series of experiments are being conducted or planned, which can enrich our understanding of the pion internal structure. Prominent among these are the JLab 12 GeV program Dudek:2012vr , the Apparatus for Meson and Baryon Experimental Research (AMBER) at CERN SPS Adams:2018pwt , the forthcoming Electron-Ion Collider (EIC) at Brookhaven National Laboratory AbdulKhalek:2021gbh , and the Electron-Ion Collider in China (EicC) Anderle:2021wcy . These initiatives are designed to probe the pion structure in various kinematic regimes. However, extracting GPDs from experimental data is fraught with challenges, including the chiral-odd nature of certain distributions, such as the transversity GPDs, and the complexity involved in pion production. On the other hand, lattice QCD results, offering complementary insights and potentially guiding experimental efforts, are highly sought after. The advent of Large Momentum Effective Theory (LaMET) Ji:2013dva ; Ji:2014gla ; Ji:2020ect in 2013 made it possible to directly calculate the Bjorken-$x$ dependence of the parton distributions, extending the scope of GPD studies beyond just the first few Mellin moments Hagler:2003jd ; Gockeler:2003jfa ; Gockeler:2005cj ; QCDSF-UKQCD:2007gdl ; LHPC:2007blg ; Alexandrou:2011nr ; Alexandrou:2013joa ; Bali:2018zgl ; Alexandrou:2019ali ; Alexandrou:2022dtc . This spurred a flurry of lattice research on nucleon and meson GPDs Chen:2019lcm ; Alexandrou:2020zbe ; Lin:2020rxa ; Lin:2021brq ; Bhattacharya:2022aob ; Bhattacharya:2023ays ; Lin:2023gxz ; Bhattacharya:2023jsc ; Holligan:2023jqh based on the quasi-distribution (see reviews in refs. Ji:2020ect ; Constantinou:2020hdm ), which is defined in terms of matrix elements of non-local operators involving quark and antiquark fields separated by a spatial distance, rather than a light-cone distance, and thus is calculable on the lattice.
@@
While there are several lattice QCD studies of the nucleon GPDs Alexandrou:2020zbe ; Lin:2020rxa ; Lin:2021brq ; Bhattacharya:2022aob ; Bhattacharya:2023ays ; Bhattacharya:2023jsc ; Holligan:2023jqh , including a few based on alternate methods CSSMQCDSFUKQCD:2021lkf ; Hannaford-Gunn:2024aix ; Dutrieux:2024umu ; Bhattacharya:2024qpp , there have been limited LaMET calculations of the pion GPDs Chen:2019lcm ; Lin:2023gxz . Traditionally, GPD calculations are conducted in the Breit (symmetric) frame, requiring the momentum transfer to be symmetrically distributed between the initial and final states. However, this approach incurs substantial computational costs. Recently, a frame-independent method was proposed to extract GPDs from lattice calculations in an arbitrary frame Bhattacharya:2022aob . This innovative approach holds the potential to significantly reduce computational expenses by working in a non-Breit (asymmetric) frame where the initial or final state momentum is fixed.
@@
Two crucial aspects of obtaining the light-cone parton distributions from the lattice calculations are the renormalization and matching process. The GPDs and quasi-GPDs (qGPDs) are typically defined in different renormalization schemes. For the qGPDs extracted from the lattice analyses, renormalization is required to eliminate the ultraviolet (UV) divergences stemming from the Wilson line. Commonly employed renormalization methods include the regularization-independent momentum-subtraction (RI/MOM) Constantinou:2017sej ; Alexandrou:2017huk ; Stewart:2017tvs ; Chen:2017mzz and various ratio schemes Radyushkin:2017cyf ; Orginos:2017kos ; Braun:2018brg ; Li:2020xml ; Fan:2020nzz . However, these methods adhere to a factorization relation with the light-cone correlation only at short distances. At long distances, they introduce nonperturbative effects Ji:2020brr , which will impact the qGPDs through the Fourier transform of the matrix elements, thereby affecting the LaMET matching results in Bjorken-$x$ space. To overcome this issue, in this work, we use a hybrid-scheme renormalization Ji:2020brr . The key point of this scheme is to renormalize the matrix elements at short and long distances separately, which can remove the linear divergence at long distances without introducing additional nonperturbative effects.
@@
Furthermore, we match the qGPDs in the lattice renormalization scheme to the light-cone GPDs in the $\overline{\rm MS}$ scheme through LaMET Ji:2015qla ; Liu:2019urm ; Ma:2022ggj ; Yao:2022vtp . During this process, the accuracy of the perturbative calculation plays a significant role in the precision of the final GPD results. In the zero-skewness limit, the matching kernel of GPDs is the same as the one for PDFs Liu:2019urm , which has been derived up to the next-to-next-to-leading order (NNLO) Li:2020xml ; Chen:2020ody . We also utilize renormalization group resummation (RGR) Gao:2021hxl ; Su:2022fiu to resum the small-$x$ logarithmic terms. Additionally, we consider leading-renormalon resummation (LRR) Holligan:2023rex ; Zhang:2023bxs , which can remove the renormalon ambiguity in the Wilson-line mass matching, to eliminate the linear power corrections Zhang:2023bxs . Therefore, we achieve a state-of-the-art calculation of the valence pion GPDs using an adapted hybrid-scheme renormalization, along with the implementation of the combined NNLO+RGR+LRR matching.
@@
In this study, we present our lattice calculations of the valence pion GPDs using the LaMET approach, featuring a very small lattice spacing of 0.04 fm within the non-Breit frame. The organization of this paper is as follows: In section 2, we review the definitions of pion GPDs and qGPDs in an arbitrary frame, outline our lattice setup, and describe the methodologies employed to extract the matrix elements from the combined analyses of the two-point and three-point correlation functions. In this section, we also detail the renormalization procedure of the qGPD matrix elements in the coordinate space using the hybrid scheme and discuss the challenges of performing the Fourier transform to get the qGPDs. Subsequently, we apply the LaMET matching approach to derive the valence light-cone GPDs in section 3. This section contains our main results, including the sensitivity of our results to the perturbative accuracy of the matching and the dependence on the renormalization scales. Finally, section 4 provides a summary of our findings.
2 Valence pion quasi-GPDs in asymmetric frame on the lattice
2.1 General considerations
@@
Our goal is to calculate the valence pion GPDs in an asymmetric frame using the frame-independent approach laid out in ref. Bhattacharya:2022aob . For this, we consider the following non-local iso-vector matrix element
@@
where ${\cal W}_{-\frac{z}{2},\frac{z}{2}}$ is the Wilson line that connects the quark and antiquark fields, ensuring the gauge invariant of the matrix element. Here, we use the following normalization of the pion states $\left\langle\pi(P)|\pi(P^{\prime})\right\rangle=E_{p}(2\pi)^{3}\delta^{3}(\mathbf{P}-\mathbf{P^{\prime}})$. Following ref. Bhattacharya:2022aob , we write the matrix element in terms of the Lorentz-invariant amplitudes $A_{i}(z\cdot\bar{P},z\cdot\Delta,\Delta^{2},z^{2})$ for $i=1,2,3$, as follows:
@@
where the skewness parameter $\xi=-\Delta^{+}/(2P^{+})$ and $t=\Delta^{2}$. However, it is possible to define the light-cone GPDs in a frame-independent, i.e., Lorentz-invariant way Bhattacharya:2022aob as
@@
In the Lorentz-invariant framework, the momentum-space GPD is defined as Bhattacharya:2022aob
@@
$\widetilde{H}(x,\xi,t,\bar{P}_{z})=\int\frac{d(z\cdot\bar{P})}{2\pi}e^{ixz\cdot\bar{P}}\widetilde{H}_{\rm LI}(z\cdot\bar{P},-2\xi(z\cdot\bar{P}),t,z^{2}).$ (9)
By studying the behavior under hermiticity and time-reversal transformations simultaneously, it was found that the amplitude $A_{3}$ is an odd function of $z\cdot\Delta=-2\xi(z\cdot\bar{P})$ Bhattacharya:2022aob ; Bhattacharya:2023jsc . In the forward limit, GPDs should smoothly approach PDFs, implying that $A_{3}$ should be equal to 0 in this work at zero skewness. In the following subsections, we will discuss our lattice setup and lattice QCD results on $A_{i}$.
2.2 Lattice setup
Our lattice calculations use the gauge ensembles provided by the HotQCD collaboration Bazavov:2019www , utilizing a 2+1 flavor setup with Highly Improved Staggered Quark (HISQ) action Follana:2006rc . The lattice configuration has dimensions of $N_{s}\times N_{t}=64^{3}\times 64$ and a lattice spacing of $a=0.04$ fm. The sea quark masses are adjusted to yield a pion mass of 160 MeV. In the valence sector, we use the Wilson-Clover action with one level of hypercubic (HYP) smearing Hasenfratz:2001hp . For the coefficient of the clover term, we utilize the tree-level tadpole-improved value computed with the fourth root of the plaquette, yielding 1.02868 Gao:2020ito . The valence quark masses in the Wilson-Clover action are tuned to $am_{q}=-0.033$, resulting in a pion mass of 300 MeV. While mixed action and partially quenched setups can theoretically introduce larger lattice artifacts compared to unitary actions, previous studies (e.g., refs. Bhattacharya:2015wna ; Mondal:2020cmt ; Gao:2020ito ; Gao:2022iex ) have shown that these effects are milder than other systematic or statistical uncertainties in practice.
Frame $t_{s}/a$ $\mathbf{n}^{f}=(n^{f}_{x},n^{f}_{y},n^{f}_{z})$ $m_{z}$ $P_{z}$[GeV] $\mathbf{n}^{\Delta}=(|n^{\Delta}_{x}|,|n^{\Delta}_{y}|,|n^{\Delta}_{z}|)$ $-t$[GeV2] #cfgs (#ex, #sl)
Breit 9,12,15,18 (1, 0, 2) 2 0.968 (2, 0, 0) 0.938 115 (1, 32)
non-Breit 9,12,15,18 (0,0,0) 0 0 (0,0,0) 0 314 (3, 96)
9,12,15,18 (0,0,2) 2 0.968 (1,2,0) 0.952 314 (4, 128)
Central to our computational approach is the use of momentum-boosted smeared sources Bali:2016lva . The quark propagators are obtained through the application of the multigrid algorithm Brannick:2007ue to invert the Wilson-Dirac operator using the QUDA software suite  Clark:2009wm ; Babich:2011np ; Clark:2016rdz on GPUs. In this work, we consider the pion boosted along the $z$-direction with momenta $P_{z}=2\pi n_{z}/(aN_{s})$ with $n_{z}$ being an integer. To obtain a good signal for the boosted pions, momentum-boosted sources and sinks are constructed using a Gaussian profile, with boost momenta $k_{z}=2\pi m_{z}/(aN_{s})$ in the $z$-direction. Source construction is done in the Coulomb gauge, and the Gaussian profile is created with the radius of 0.208 fm Izubuchi:2019lyk ; Gao:2020ito . Employing these quark propagators, we have computed both the two-point and three-point hadron correlation functions. To increase statistics per configuration, we combined multiple exact (high-precision) and sloppy (low-precision) samples and implemented the all-mode averaging (AMA) technique Shintani:2014vja . The lattice parameters used in this study are summarized in table 1.
@@ 2.2 Lattice setup
The approach used to obtain the pion matrix element here closely follows our previous works Gao:2020ito ; Gao:2021dbh ; Gao:2021xsm ; Gao:2022iex ; Gao:2022vyh ; Ding:2024lfj . We fit the large $t_{s}$ behavior of the two-point correlation functions with two or three states and obtain the energies of these states for different momenta. The corresponding energy levels are then used in the analyses of the three-point function and the extraction of the pion matrix elements $M^{\mu}$. We consider the following ratio of
$R^{\mu}(\mathbf{P}^{f},\mathbf{P}^{i};t_{s},\tau;z)$ $\equiv\sqrt{E_{0}^{f}E_{0}^{i}}\frac{C_{{\rm 3pt}}^{\mu}(\mathbf{P}^{f},\mathbf{P^{i}};t_{s},\tau;z)}{C_{{\rm 2pt}}(\mathbf{P}^{f},t_{s})}$
×$\qquad\times\left[\frac{C_{{\rm 2pt}}(\mathbf{P}^{i},t_{s}-\tau)C_{{\rm 2pt}}(\mathbf{P}^{f},\tau)C_{{\rm 2pt}}(\mathbf{P}^{f},t_{s})}{C_{{\rm 2pt}}(\mathbf{P}^{f},t_{s}-\tau)C_{{\rm 2pt}}(\mathbf{P}^{i},\tau)C_{{\rm 2pt}}(\mathbf{P}^{i},t_{s})}\right]^{1/2},$ (12)
where $E_{0}^{i}$ and $E_{0}^{f}$ correspond to the ground-state pion energies in the initial and final states, respectively, and the kinematic factor $\sqrt{E_{0}^{f}E_{0}^{i}}$ is used for the normalization. For large $t_{s}$ and $\tau$, this ratio gives the matrix element $M^{\mu}$. We assume that for large $t_{s}$ and $\tau$, two or three states mainly contribute to $R^{\mu}(\mathbf{P}^{f},\mathbf{P}^{i};t_{s},\tau;z)$ and perform the corresponding fits to obtain the matrix element. More details of the fits are provided in appendix A.
2.3 Lattice results on the Lorentz-invariant amplitudes